OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 2(5y - 1) + 4y = -4 2(-5y + 1) + 4y = -4 2x + 4(-5y + 1) = -4 2x + 4(5y - 1) = -4

OpenStudy (anonymous):

a 2(5y - 1) + 4y = -4 b 2(-5y + 1) + 4y = -4 c 2x + 4(-5y + 1) = -4 d 2x + 4(5y - 1) = -4

OpenStudy (mathmale):

Futuristic: Please solve the first equation, -x + 5y = 1, for x. Then substitute your new formula for x into the second equation, replacing x in the second equation with your formula for x.